OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (−5, 1). y = −5 x = 1 y = 1 x = −5 i will give medal

OpenStudy (alexandervonhumboldt2):

the equation has form y=mx+b where m is slope and b is y-intercept. Slope of horizontal lines is 0. How substitute one of our points into the equation and solve 1=-5*0+b b=1. thus the equation is y=1:

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@AlexandervonHumboldt2 how do i give u a medal

OpenStudy (alexandervonhumboldt2):

by clicking best responce